# STORM: A Spatio-Temporal Factor Model Based on Dual Vector Quantized Variational Autoencoders for Financial Trading

### v2

@@ Abstract.
In financial trading, factor models are widely used to price assets and capture excess returns from mispricing. Recently, we have witnessed the rise of variational autoencoder-based latent factor models, which learn latent factors self-adaptively. While these models focus on modeling overall market conditions, they often fail to effectively capture the temporal patterns of individual stocks. Additionally, representing multiple factors as single values simplifies the model but limits its ability to capture complex relationships and dependencies. As a result, the learned factors are of low quality and lack diversity, reducing their effectiveness and robustness across different trading periods. To address these issues, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders, named STORM, which extracts features of stocks from temporal and spatial perspectives, then fuses and aligns these features at the fine-grained and semantic level, and represents the factors as multi-dimensional embeddings. The discrete codebooks cluster similar factor embeddings, ensuring orthogonality and diversity, which helps distinguish between different factors and enables factor selection in financial trading. To show the performance of the proposed factor model, we apply it to two downstream experiments: portfolio management on two stock datasets and individual trading tasks on six specific stocks. The extensive experiments demonstrate STORM’s flexibility in adapting to downstream tasks and superior performance over baseline models. Code is available in PyTorch [^1] https://github.com/DVampire/Storm.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Proceedings of the ACM Web Conference 2025; April 28–May 02, 2025; Sydney, Australiaisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Data miningccs: Computing methodologies Machine learningccs: Applied computing Electronic commerce
@@ 1. Introduction
In order to address the challenges, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders (VQ-VAE), named STORM, with the architecture shown in Figure 1. We represent the factors as multi-dimensional embeddings to capture the complexity and nonlinearity of financial data, thereby improving the factors’ ability to account for factor returns (CH1). Additionally, we develop a dual VQ-VAE architecture to capture cross-sectional and time-series features, considering both spatial and temporal perspectives [^2] We treat cross-sectional as spatial and time-series as temporal. These terms will be used interchangeably unless specified.. We integrate cross-sectional and time-series features at both fine-grained and semantic levels to construct effective factors. To further enhance the model, we introduce diversity loss and orthogonality loss to ensure factor embedding diversity and independence (CH2). Furthermore, we treat the embeddings in the codebook as cluster centers for factor embeddings, effectively using them as class tokens to distinguish between factors. This approach makes the factor differentiation and selection process clear and transparent (CH3). Specifically, our contributions are three-fold:
  • We design a dual VQ-VAE architecture to construct cross-sectional and time-series factors from both spatial and temporal perspectives.
@@
  • Experiments on two U.S. stock markets across the stock future return prediction task and two downstream tasks (i.e., portfolio management and algorithmic trading) show that STORM outperforms all baselines across 2 prediction metrics and 6 standard financial metrics.
@@ 2.1. Factor Model
Factor model has attracted extensive research focus from the finance community. Financial experts have carefully selected a wide range of factors from macroeconomic (Sharpe, 1964), market changes (Fama and French, 1992), individual stock fundamentals (Greenblatt, 2010), and other perspectives to price assets in an attempt to capture excess returns from mispricing. Recently, latent factor models have emerged that improve investment returns by studying latent factors self-adaptively. Specifically, (Gu et al., 2021) learns latent factors depend on assets characteristics, (Duan et al., 2022) learns cross-sectional factors by predicting future returns, and (Wei et al., 2023) learns market-augmented latent factors. However, these VAE-based latent factor models may suffer from a lack of robustness in representing low signal-to-noise ratio markets, leading to reduced reconstruction and generation capabilities. Additionally, their effectiveness is constrained by a primary focus on with insufficient attention to the temporal perspective.
@@ 2.2. Financial Trading
Financial trading has attracted significant attention from both the finance and AI communities. Quantitative traders leverage mathematical models and algorithms to automatically identify trading opportunities (Sun et al., 2024). Within the Fin-tech field, research tasks include portfolio management (PM), algorithmic trading (AT), order execution (Fang et al., 2021), market making (Spooner and Savani, 2020), etc.
PM focuses on optimizing wealth allocation across assets (Li and Hoi, 2014). Its development has evolved from simple rule-based methods (Poterba and Summers, 1988) to prediction-based approaches using machine learning (Ke et al., 2017) and deep learning (Qin et al., 2017; Zhao et al., 2022) for predicting returns or price movements. Despite recent integration of reinforcement learning (RL) in PM (Zhang et al., 2024b), the latent factor models still rely on deep learning for market modeling and precise predictions.
AT involves trading financial assets using algorithm-generated signals. Similar to PM, there are rule-based and supervised learning methods (Yang et al., 2020) in the AT task. In recent years, RL-based methods (Sun et al., 2022) have become dominant for their superiority in complex sequential decision-making.
@@ 2.3. Vector Quantized Variational Autoencoder
VQ-VAE has made significant contributions across several areas of deep learning, particularly in representation learning, generative modeling, and time series analysis. Initially introduced by Van Den Oord et al., VQ-VAE has showcased its exceptional capabilities in processing high-dimensional data, e.g., images (Razavi et al., 2019), audio (Zhang et al., 2024a), and video (Yan et al., 2021), with a strong emphasis on learning discrete latent representations. Moreover, VQ-VAE has proven highly effective in handling complex time-dependent data (Talukder et al., 2024), advancing the development of robust generative models specifically in the context of temporal analysis.
@@ 3.1. Factor Model
$\mathbb{E}[\cdot]$ denotes the expectation operator. $\mathbb{E}[R_{i}^{e}]$ is the expected excess return. $\beta_{i}$ is factor exposure (or factor loading) of asset $i$, and $\lambda$ is factor expected return (or factor risk premium). $\alpha_{i}$ is pricing error.
According to (Duan et al., 2022), they define the factor model in the spatial perspective, which only focuses on why the expected returns of different assets vary, rather than how the returns of each asset change over time. Therefore, we expand it from the temporal perspective and introduce the general functional form of the latent factor model:
@@
where $y_{i,t}$ denotes the future return of stock $i$ at timestep $t$, $\mathbf{z}_{i,t}$ denotes $K$ factor returns, and $\epsilon_{i,t}$ is the noise. From the spatial, Equation 2 can be expressed as $\mathbf{y}_{i}=\bm{\alpha}_{i}+\sum_{k=1}^{K}\bm{\beta}_{i}^{k}\mathbf{z}_{i}^{k}+\bm{\epsilon}_{i}$, which we regard as the cross-sectional factor model. From the temporal perspective, Equation 2 can be expressed as $\mathbf{y}_{t}=\bm{\alpha}_{t}+\sum_{k=1}^{K}\bm{\beta}_{t}^{k}\mathbf{z}_{t}^{k}+\bm{\epsilon}_{t}$, which we regard as the time-series factor model. The cross-sectional factor model focuses on the differences in returns of different assets at the same point in time, while the time-series factor model focuses on the changes in returns of a single asset at different points in time.
@@ 3.2. Problem Formulation
To assess the effectiveness and generating ability of factors, we predict and evaluate the stock future returns. To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (Wang et al., 2021; Zhang et al., 2024b; Ye et al., 2020) and algorithmic trading (Sun et al., 2022; Liu et al., 2020), to showcase the investment capabilities.
Observed Data: we utilize the stock’s historical price data $\mathbf{p}$(i.e., open, high, low, and close) and technical indicators $\mathbf{d}$ as observed variables $\mathbf{x}:=[\mathbf{p},\mathbf{d}]\in\mathbb{R}^{N\times W\times D}$, within a window size of $W$. The historical price data of stock $i$ at time $t$ is denoted as $\mathbf{p}_{i,t}:=[p_{i,t}^{o},p_{i,t}^{h},p_{i,t}^{l},p_{i,t}^{c}]\in\mathbb{R}^{D_{1}}$. The technical indicators which are calculated from price and volume values, are denoted as $\mathbf{d}_{i,t}\in\mathbb{R}^{D_{2}}$. Each stock is represented by $D=D_{1}+D_{2}$ features per trading day, which we then use to predict the stock’s future returns, defined as $\mathbf{y}_{i,t+1}:=\frac{p_{i,t+1}-p_{i,t}}{p_{i,t}}$.
Downstream tasks:
  • Portfolio management task aims to construct a portfolio based on the predicted returns to test the profitability of the factor model. At each trading day $t$, we obtain the portfolio weight $\mathbf{w}_{t}=[w_{t}^{0},w_{t}^{1},...,w_{t}^{N-1}]\in\mathbb{R}^{N}$ for $N$ stocks. Each stock $i$ is assigned a weight $w_{t}^{i}$ representing its portfolio proportion, subject to the constraint that $\sum_{i=0}^{N-1}w_{t}^{i}=1$ for full investment.
@@ 4.1. Dual VQ-VAE Structure
The dual VQ-VAE architecture processes stock data from spatial and temporal perspectives by patching the data in both modules. Then, the encoder extracts time-series (TS) and cross-sectional (CS) features, which are subsequently clustered using a learnable codebook for quantized representation. Finally, the decoder reconstructs the data, capturing complex patterns and enhancing forecasting performance.
@@ 4.1.2. TS and CS Encoders
Considering the Transformer’s (Vaswani et al., 2017) remarkable ability to capture both local and global dependencies via self-attention and patching strategies, it is particularly well-suited for long-term time-series forecasting (Nie et al., 2022). Therefore, we employ stacked transformer blocks as the encoders in the TS and CS modules to capture complex time-series and cross-sectional patterns, ultimately improving the accuracy and robustness of feature extraction.
@@ 4.1.3. Codebook Construction and Optimization
The VQ-VAE offers significant advantages in the latent factor model, primarily through its use of discrete codebook embeddings: i) Discretization Benefits: The discretization process inherent in VQ-VAE helps in clustering distinct and meaningful factors, providing orthogonality and diversity which improves upon the factors derived from traditional methods like FactorVAE. ii) Explicit Factor Selection: The discrete token indices facilitate an explicit factor selection process, identifying the most relevant factors that influence stock returns and thus enhancing prediction accuracy. iii) Noise Reduction: The factor selection process filters out irrelevant or redundant information, reducing noise and improving the overall robustness of the model.
@@
(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}\vspace{-0.2cm}$
@@
However, the lack of independence among factors can lead to multicollinearity, causing the model to overfit and are less robust to various market situations. To address this issue, we impose an orthogonality constraint on the codebook, as recent studies (Shin et al., 2023) have shown that enforcing orthogonality allows discretized codes to maintain translation equivariance. The orthogonality loss is:
@@ 4.2. TS and CS Decoders
Finally, the encoders, decoders, and codebooks are jointly learned by minimizing the following loss objectives:
@@ 4.3. Factor Module
The factor module aims to integrate the TS and CS latent features through feature fusion and alignment. Then, through prior-posterior learning to generate factor embeddings from latent features to predict stock future returns.
@@ 4.3.1. Feature Fusion and Alignment
As mentioned above we employ different patching methods to ensure that both time-series and cross-sectional features of the data were accurately captured. However, these methods result in two parallel processes that do not align at the feature level. To address this limitation, we employ a cross-attention mechanism to enhance the interaction and fusion of features at a fine-grained level, improving the model’s dynamic understanding of the input data. Additionally, contrastive learning is used at the semantic level to enhance feature representation by emphasizing semantic similarities and differences.
Cross-Attention Mechanism. In order to fuse the features in the TS and CS modules at a fine-grained level, we leverage a multi-scale encoder to combine information between two patches and produce stronger fusion features for factor generations, motivated by CrossViT (Chen et al., 2021). The encoder contains cross-attention layers that achieve both accuracy and efficiency.
Contrastive Learning. We incorporate contrastive learning, inspired by its success in aligning text and image features through a shared embedding space (Li et al., 2023; Radford et al., 2021), to enhance semantic consistency between time-series and cross-sectional features. In the factor module, contrastive loss ensures that similar features cluster together while dissimilar features are pushed apart in the embedding space.

### v3

@@ Abstract.
In financial trading, factor models are widely used to price assets and capture excess returns from mispricing. Recently, we have witnessed the rise of variational autoencoder-based latent factor models, which learn latent factors self-adaptively. While these models focus on modeling overall market conditions, they often fail to effectively capture the temporal patterns of individual stocks. Additionally, representing multiple factors as single values simplifies the model but limits its ability to capture complex relationships and dependencies. As a result, the learned factors are of low quality and lack diversity, reducing their effectiveness and robustness across different trading periods. To address these issues, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders, named STORM, which extracts features of stocks from temporal and spatial perspectives, then fuses and aligns these features at the fine-grained and semantic level, and represents the factors as multi-dimensional embeddings. The discrete codebooks cluster similar factor embeddings, ensuring orthogonality and diversity, which helps distinguish between different factors and enables factor selection in financial trading. To show the performance of the proposed factor model, we apply it to two downstream experiments: portfolio management on two stock datasets and individual trading tasks on six specific stocks. The extensive experiments demonstrate STORM’s flexibility in adapting to downstream tasks and superior performance over baseline models. Code is available in PyTorch [^1] https://anonymous.4open.science/r/Storm-2823.
@@
conference: Proceedings of the 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining 2025; August 03–August 07, 2025; Toronto, Canada
@@ 1. Introduction
In order to address the challenges, we propose a Spatio-Temporal factOR Model based on dual vector quantized variational autoencoders (VQ-VAE), named STORM, with the architecture shown in Figure 1. The model encodes factors as multi-dimensional embeddings, effectively capturing the complexity and nonlinearity inherent in financial data, and enhancing the capacity of factors to account for factor returns (CH1). Additionally, we develop a dual VQ-VAE architecture to capture cross-sectional and time-series features, considering both spatial and temporal perspectives [^2] We treat cross-sectional as spatial and time-series as temporal. These terms will be used interchangeably unless specified.. By integrating these features at both fine-grained and semantic levels, the model constructs more effective factors. Through diversity and orthogonality loss constraints, we ensure representation independence among learned representations (CH2). Furthermore, codebook embeddings act as cluster centers, serving as class tokens to categorize factor embeddings. This strategy provides clarity and transparency to the differentiation and selection process of factors (CH3). Specifically, our contributions are three-fold:
  • We design a dual VQ-VAE architecture that systematically constructs cross-sectional and time-series factors by capturing both spatial and temporal dependencies.
@@
  • We develop STORM as a versatile all-rounder capable of supporting three tasks simultaneously, whereas other methods tend to be specialized in specific tasks.
  • Experiments on seven datasets from two U.S. stock markets, covering the stock future return prediction task (DJ30 and SP500), portfolio management (DJ30 and SP500), and algorithmic trading (five stocks), show that STORM outperforms all baselines across two prediction metrics and six standard financial metrics.
@@ 2.1. Factor Model
Factor model has attracted extensive research focus from the finance community. Financial experts have carefully selected a wide range of factors from macroeconomic (Sharpe, 1964), market changes (Fama and French, 1992), individual stock fundamentals (Greenblatt, 2010), and other perspectives to price assets in an attempt to capture excess returns from mispricing. Recently, latent factor models have emerged that improve investment returns by studying latent factors self-adaptively. Specifically, (Gu et al., 2021) learns latent factors depend on asset characteristics, (Duan et al., 2022) extracts cross-sectional factors by predicting future returns, and (Wei et al., 2023) learns market-augmented latent factors. However, these VAE-based latent factor models may suffer from a lack of robustness in representing low signal-to-noise ratio markets, leading to reduced reconstruction and generation capabilities. Furthermore, their effectiveness is constrained by an insufficient focus on the temporal perspective.
@@ 2.2. Financial Trading
Financial trading has attracted significant attention from both the finance and AI communities. Quantitative traders leverage mathematical models and algorithms to automatically identify trading opportunities (Sun et al., 2024). Within the FinTech domain, key research areas encompass portfolio management (PM), algorithmic trading (AT), order execution (Fang et al., 2021), market making (Spooner and Savani, 2020), etc. In this work, we focus on PM and AT, as they are fundamental to developing intelligent financial decision-making systems.
PM focuses on optimizing wealth allocation across assets (Li and Hoi, 2014). Its development has evolved from simple rule-based methods (Poterba and Summers, 1988) to prediction-based approaches using machine learning (Ke et al., 2017) and deep learning (Qin et al., 2017; Zhao et al., 2022) for predicting returns or price movements. Despite recent integration of reinforcement learning (RL) in PM (Zhang et al., 2024b), the latent factor models still predominantly rely on deep learning for market modeling and precise predictions.
AT, on the other hand, involves executing trades with algorithm-generated signals. Like PM, AT has seen a transition from rule-based and supervised learning techniques (Yang et al., 2020) to RL-driven approaches. In particular, RL-based methods (Sun et al., 2022) have gained prominence due to their ability to navigate complex sequential decision-making problems, further enhancing trade execution efficiency and adaptability in dynamic market environments.
@@ 2.3. Vector Quantized Variational Autoencoder
Vector Quantized Variational Autoencoder (VQ-VAE) has emerged as a fundamental framework in deep learning research, establishing methodological significance across core domains: representation learning, generative modeling, and time series analysis. Initially introduced by Van Den Oord et al., VQ-VAE has showcased its capabilities in processing high-dimensional data across diverse modalities, including images (Razavi et al., 2019), audio (Zhang et al., 2024a), and video (Yan et al., 2021), with a strong emphasis on learning discrete latent representations. Particularly noteworthy is the model’s demonstrated proficiency in capturing temporal dependencies, where recent studies like (Talukder et al., 2024) have validated its superior performance in modeling sequential patterns compared to conventional continuous latent space approaches, thereby advancing the development of robust generative architectures specifically optimized for temporal dynamics analysis.
@@ 3.1. Factor Model
Here, $\mathbb{E}[\cdot]$ denotes the expectation operator, and $\mathbb{E}[R_{i}^{e}]$ represents the expected excess return of asset $i$. $\beta_{i}$ is factor exposure (or factor loading) of asset $i$, and $\lambda$ is factor expected return (or factor risk premium). Lastly, $\alpha_{i}$ captures pricing error, reflecting any deviations from the model’s predicted returns.
According to (Duan et al., 2022), the factor model is formulated from a spatial perspective, primarily examining cross-sectional variations in expected returns across different assets, rather than capturing the temporal evolution of individual asset returns. Therefore, we expand it from the temporal perspective and introduce the general functional form of the latent factor model:
@@
where $y_{i,t}$ denotes the future return of stock $i$ at timestep $t$, $\mathbf{z}_{i,t}$ denotes $K$ factor returns, and $\epsilon_{i,t}$ is the noise. From the spatial, Equation 2 can be expressed as $\mathbf{y}_{i}=\bm{\alpha}_{i}+\sum_{k=1}^{K}\bm{\beta}_{i}^{k}\mathbf{z}_{i}^{k}+\bm{\epsilon}_{i}$, which we regard as the cross-sectional factor model. From the temporal perspective, Equation 2 can be expressed as $\mathbf{y}_{t}=\bm{\alpha}_{t}+\sum_{k=1}^{K}\bm{\beta}_{t}^{k}\mathbf{z}_{t}^{k}+\bm{\epsilon}_{t}$, which we regard as the time-series factor model. The cross-sectional factor model focuses on the differences in returns of different assets at the same point in time, while the time-series factor model focuses on the changes in returns of a single asset at different time points.
@@ 3.2. Problem Formulation
To assess the effectiveness and generating ability of factors, we predict and evaluate the stock future returns. To show the adaptability of the latent factor model with multiple downstream tasks, we employ two financial trading tasks, i.e., portfolio management (Wang et al., 2021a; Zhang et al., 2024b; Ye et al., 2020) and algorithmic trading (Sun et al., 2022; Liu et al., 2020), to showcase the model’s investment capabilities.
Observed Data. We utilize the stock’s historical price data $\mathbf{p}$ (i.e., open, high, low, and close) and technical indicators $\mathbf{d}$ as observed variables $\mathbf{x}:=[\mathbf{p},\mathbf{d}]\in\mathbb{R}^{N\times W\times D}$, within a window size of $W$. The historical price data of stock $i$ at time $t$ is denoted as $\mathbf{p}_{i,t}:=[p_{i,t}^{o},p_{i,t}^{h},p_{i,t}^{l},p_{i,t}^{c}]\in\mathbb{R}^{D_{1}}$. The technical indicators which are calculated from price and volume values, are denoted as $\mathbf{d}_{i,t}\in\mathbb{R}^{D_{2}}$. Each stock is represented by $D=D_{1}+D_{2}$ features per trading day, which we then use to predict the stock’s future returns, defined as $\mathbf{y}_{i,t+1}:=\frac{p_{i,t+1}-p_{i,t}}{p_{i,t}}$.
Downstream tasks. These primarily encompass the two most widely studied areas in quantitative finance – Portfolio management and Algorithmic trading:
  • Portfolio management task aims to construct an optimal portfolio based on the predicted returns to test the profitability of the factor model. At each trading day $t$, we obtain the portfolio weight $\mathbf{w}_{t}=[w_{t}^{0},w_{t}^{1},...,w_{t}^{N-1}]\in\mathbb{R}^{N}$ for $N$ stocks. Each stock $i$ is assigned a weight $w_{t}^{i}$ representing its portfolio proportion within the portfolio. To ensure full capital deployment, the allocation must satisfy the budge constraint $\sum_{i=0}^{N-1}w_{t}^{i}=1$.
@@ 4.1. Dual VQ-VAE Structure
The dual VQ-VAE architecture processes stock data from both spatial and temporal perspectives by patching the data within each module. The encoder then extracts time-series (TS) and cross-sectional (CS) features, which are subsequently clustered using a learnable codebook for quantized representation. Finally, the decoder reconstructs the data, effectively capturing complex patterns and enhancing model’s forecasting performance.
@@ 4.1.2. TS and CS Encoders
Considering the Transformer’s (Vaswani et al., 2017) remarkable ability to capture both local and global dependencies via self-attention and patching strategies, it is particularly well-suited for long-term time-series forecasting (Nie et al., 2022). Therefore, we employ stacked transformer blocks as the encoders in both the TS and CS modules to capture complex time-series and cross-sectional patterns. This architecture enhances the accuracy and robustness of feature extraction, ultimately improving forecasting performance.
While transformers are widely applicable across domains such as natural language processing (NLP) and computer vision (CV), STORM’s flexible encoder-decoder framework allows for the integration of alternative architectures, such as long short-term memory (LSTM) (Wang et al., 2021b) and temporal convolutional networks (TCN) (Yu and Koltun, 2015).
@@ 4.1.3. Codebook Construction and Optimization
The VQ-VAE offers significant advantages in the latent factor model, primarily through its use of discrete codebook embeddings: i) Discretization Benefits: The discretization process inherent in VQ-VAE helps in clustering distinct and meaningful factors, providing orthogonality and diversity which improves upon the factors derived from traditional methods like FactorVAE. ii) Explicit Factor Selection: The discrete token indices facilitate an explicit factor selection process, identifying the most relevant factors that influence stock returns and thus enhancing prediction accuracy. iii) Noise Reduction: By filtering out irrelevant or redundant information during factor selection, VQ-VAE mitigates noise, contributing to improved model robustness and more reliable financial insights.
@@
(4) $\mathcal{L}_{\text{div}}=\frac{1}{GK}\sum_{g=1}^{G}\sum_{k=1}^{K}\bar{p}_{g,k}\log{\bar{p}_{g,k}}\text{,}\vspace{-0.1cm}$
@@
However, the lack of independence among factors can introduce multicollinearity, making the model prone to overfitting and reducing its robustness across different market conditions. To address this issue, we impose an orthogonality constraint on the codebook, as recent studies (Shin et al., 2023) have shown that enforcing orthogonality allows discretized codes to maintain translation equivariance. The orthogonality loss is:
@@ 4.2. TS and CS Decoders
Finally, the encoders, decoders, and codebooks are jointly optimized by minimizing the following loss objectives:
@@ 4.3. Factor Module
The factor module aims to integrate the TS and CS latent features through feature fusion and alignment. Afterward, it utilizes prior-posterior learning to generate factor embeddings from latent features to predict stock future returns.
@@ 4.3.1. Feature Fusion and Alignment
As mentioned above we employ different patching methods to ensure that both time-series and cross-sectional features of the data were accurately captured. However, these methods result in two parallel processes that do not align at the feature level. To address the limitation, we employ a cross-attention mechanism to enhance the interaction and fusion of features at a fine-grained level, improving the model’s dynamic understanding of the input data. Additionally, contrastive learning is used at the semantic level to enhance feature representation by emphasizing semantic similarities and differences.
Cross-Attention Mechanism. In order to fuse the features in the TS and CS modules at a fine-grained level, we leverage a multi-scale encoder to combine information between two patches and produce stronger fusion features for factor generations, motivated by CrossViT (Chen et al., 2021). The architecture incorporates cross-attention layers, balancing both accuracy and computational efficiency, ensuring that factors derived from the combined features are more representative of market dynamics.
Contrastive Learning. To further enhance semantic consistency between TS and CS features, we integrate contrastive learning, drawing inspiration from its success in aligning multimodal representations (Li et al., 2023; Radford et al., 2021). Within the factor module, a contrastive loss encourages similar features to cluster closely while pushing apart dissimilar ones in the embedding space.